OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 2 -2 1 -1

OpenStudy (anonymous):

well you could use substitution, or addition/subtraction of equations. your choice really

OpenStudy (anonymous):

2x + 3y = 4 2x - 4y = -10 7y = 14 y = 2 plug in y into one of your original equations 2x + 6 = 4 2x = -2 x = -1

OpenStudy (anonymous):

that is using subtraction of equations

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw